Consider the following recent financials for XYZ Corporation: Income Statement Balance Sheet Sales 72,429Assets 178,884 Debt 44,955 Costs 43,457 Equity 133,929 EBIT 28,972 Taxes @ 38% 11,009 Total 178,884Total 178,884 Net income|17963 Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2.57

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2,577 was paid, and the company wishes to maintain a constant payout ratio. Next years sales are projected to grow by 24%. What is the pro-forma value for equity? (Round answer to 2 decimal places. Do not round intermediate calculations. Also, do not calculate the numbers given in the income statement and balance sheet, such as the Taxes and Net income. Take them as given.). What is the external financing needed using the pro-forma approach? (Round answer to 2 decimal places. Do not round intermediate calculations. Also, do not calculate the numbers given in the income statement and balance sheet, such as the Taxes and Net income. Take them as given.). What is the internal growth rate? (Report answer in percentage terms and round to 2 decimal places. Do not round intermediate calculations). What is the sustainable growth rate? (Report answer in percentage terms and round to 2 decimal places. Do not round intermediate calculations). Tonic: Financial Models (EFN & Growthl
Answer #1

Calculation of proforma value of equity:

Given, Sales = $72,429; Costs = $43,457; Assets = $178,884

It was given that costs and assets are proportional to sales. Debt and equity are not.

Costs to sales ratio = $43,457/ $72,429 = 0.5999

Assets to sales ratio = $178,884/ $72,429 = 2.4697

If next year sales grow by 24% then sales = $72,429 * 124% = $89,811.96

Next year costs = next year sales * costs to sales ratio

= $89,811.96 * 0.5999 = $53,878.194

Next year assets = next year sales * assets to sales ratio

= $89,811.96 * 2.4697 = $221,808.597

next year EBIT = next year sales - next year costs = $89,811.96 - $53,878.194 = $35,933.766

next year net income = next year EBIT - tax = $35,933.766 - (30% * $35,933.766) = $25,153.6362

Pro forma value of equity = assets + next year net income = $178,884 + $25,153.6362 = $204,037.64

Calculation of net financing needed:

Net financing needed = next year assets - proforma value of equity = $221,808.597 - $204,037.64 = $17,770.96

Internal Growth rate (IGR):

Net income = $17,963

Dividend paid = $2,577

Dividend pay-out ratio = ($2,577/ $17,963) * 100 = 14.35%

assets = $178,884

return of assets = net income/ assets = ($17,963/ $178,884) * 100 = 10.04%

Internal growth rate = (1 - dividend payout ratio) * return on assets = (1 - 14.35%) * 10.04%

= (0.8565) * 10.04% = 0.0859 = 8.59%

IGR = 8.59%

Sustainable growth rate (SGR):

net income = $17,963

Dividend pay-out ratio = 14.35%

equity = $133,929

return on equity = net income/ equity = $17,963/ $133,929 = 0.13412 = 13.41%

SGR = (1 - dividend payout ratio) * return on equity = (1 - 14.35%) * 13.41%

= (0.8565) * 13.41% = 0.11485 = 11.49%

SGR = 11.49%